For the function f, construct and simplify the difference quotient
f(x + h)-f(x)
_____________
h
f(x)=7x^2 +3x

Answer by Boreal(15235)

f(x+h)=7(x^2+2xh+h^2)+3(x+h)=7x^2+14xh+3x+3h+h^2)
f(x)=7x^2+3x
subtract and have the numerator of 14xh+3h+h^2
divide by h
and have 14x+3+h
(and as h goes to 0 have 14x+3, which is the derivative)